OpenStudy (anonymous):

Simplify: \[\sqrt{c}(3+2\sqrt{5})\]

OpenStudy (anonymous):

foil

OpenStudy (anonymous):

multiply by \[\sqrt{c}\]

OpenStudy (anonymous):

So far I have \[3\sqrt{c}\times \sqrt{5c}\] is that right?

OpenStudy (anonymous):

I don't know what to do next though.

OpenStudy (anonymous):

oops forgot the two in front of sqrt(5c)

OpenStudy (anonymous):

yes fixing the 2 does it,it will be correct